OpenStudy (anonymous):

A rocket is launched from atop a 65-foot cliff with an initial velocity of 113 ft/s. The height of the rocket above the ground at time t is given by -16t^2+141t+115 When will the rocket hit the ground after it is launched? Round to the nearest tenth of a second.

OpenStudy (campbell_st):

it will hit the ground when h(t) = 0... or it has zero height.. so you need to solve \[-16t^2 + 141t + 115 = 0\] the general quadratic formula ma come in handy

OpenStudy (anonymous):

Which is?

OpenStudy (anonymous):

Im just very confused on the whole section. @campbell_st

OpenStudy (campbell_st):

\[t = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you have a = -16, b = 141 and c = 115 substitute the values and calculate t. only use the positive answer.

OpenStudy (anonymous):

thanks :)